OpenStudy (anonymous):

Question about 'The Sum of; 'Why is Step2 and Step3 equal to each other?

OpenStudy (anonymous):

Okay?

OpenStudy (anonymous):

You don't get this one?\[ \sum_{\Omega=1}^n\Omega =\frac{n(n+1)}{2} \]

OpenStudy (anonymous):

\[ \sum_{\Omega = 1}^{n} 1 =\underbrace{ 1+1+\ldots +1}_n = n \]

OpenStudy (anonymous):

Is the first one equal to each other? why is that?

OpenStudy (anonymous):

is that a definition?

OpenStudy (anonymous):

axiom? :/

OpenStudy (anonymous):

So you want me to show you why this is true\[ \sum_{\Omega=1}^n\Omega =\frac{n(n+1)}{2} \]is that correct?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

First of all \[ \sum_{\Omega=1}^n n-(\Omega-1) = n + (n-1) + \ldots + 1 = \sum_{\Omega=1}^n \Omega \]Does that make sense?

OpenStudy (anonymous):

Think about it \(\Omega = 1 \implies n-(\Omega - 1) = n\), and \(\Omega = n \implies n-(\Omega - 1) = 1\). Basically it's the same sum, but with the terms in the same direction.

OpenStudy (anonymous):

terms in opposite direction.

OpenStudy (anonymous):

Ah, thank you I get it now.